9/7/99 Day Shift Report

Richard_Coleman (coleman@fsgi03)
Tue, 07 Sep 1999 15:31:40 CDT

Shift Summary for day shift, 7/Sep/99

On Shift: VP,RC

Hours of beam available at all intensities: 7.75
Hours of beam available at nominal intensity: 7.75
Hours of E832 data (special runs not included): 7.5
Hours of special runs (laser scans, muon runs...): 0.25
Accounting of lost time: NM2S3 power supply water leak- 0.5 hrs
Re-starting run after init tests - 0.25 hrs
Aborts due to TeV kicker - 0.25 hrs

========================================================================
Runs ended this shift: 14452,14453,14456

14452 6.5 M events to tape with 9.3K of vac00 off-line

Tapes brought to FCC this shift: none
(including tapes for online split)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift: Smooth data taking for the most part

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14452 8:50 NM2S3 power supply off, water leak

14453 9:00 End Run 14452, Started Muon Run

14453 9:18 updated DCGM constants at end of run

9:20 tested init for BEAM_E832_nothresh
and BEAM_E832_regout

14456 9:45 Lowered intensity request from 7 to 4 E12 as per Ed
Beam returns
Outbox problems due to previous init of non-standard
configurations- Data not usable

14457 9:55 Drt_fresh_start and start new run

14457 14:00 Many beam aborts ~50% for last 0.5 hrs
MCR called aborts were adjustments to TeV abort
kicker which had caused two quenches on previous
shifts.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift